can u help me with this equation: -3(h-6)=5(2h+3)
Answer by yhtong(11)

ok, heres what you do,
first solve all that you can solve. in this case its:
-3(h-6)=5(2h+3)
-3h+18=10h+15
then, bring all values that has the unknown for instance, -3h and 10h to the left hand side (LHS) and bring all the values without unknown to the right hand side (RHS) . Since -3h is already in the LHS, therefore you dont need to do anything to it. Its the same for 15 that is already in the RHS. That makes,
-3h-10h=15-18 (remember to change their sign from + to - or vice versa when
bringing them to the opposite side)
then, you factor the ones in the LHS
h(-3-10)=15-18
Since the unknown can be seperated with the coeficient, bring the coeficient, (-3-10) to the RHS but before that solve it first, solve the equation in the RHS as well
h(-13)=-3
h =-3/-13
h =3/13
therefore your answer is 3/13